OpenStudy (anonymous):

Are 2 times x plus 4 all over 3 times x plus 6 and 2 over 3 equivalent expressions? Using complete sentences explain why or why not.

OpenStudy (unklerhaukus):

\[\frac{2\times x +4 }{3\times x+6}\]

OpenStudy (unklerhaukus):

is that right?

OpenStudy (anonymous):

yes

OpenStudy (unklerhaukus):

can you factorise the numerator and the denominator

OpenStudy (anonymous):

its ok I figured out the question but thanks any way

OpenStudy (unklerhaukus):

what did you find?

OpenStudy (anonymous):

that yes it is a equivalent expression.

OpenStudy (unklerhaukus):

did you form a complete sentence ?

OpenStudy (anonymous):

This is because when you factorize those two factors, you are able to get 2/3 with the same x+2 Cancel the x+2 since they are 1. (Egs. 1/1 is equal to 1. 2/2 is equal to 1. So (x+2)/(x+2) is also equals to 1) You get 2/3 as the answer.

OpenStudy (unklerhaukus):

Well thats the right method, but you should say "..factorize the numerator and denominator" or "..factorize the the top and bottom of the fraction" The denominator isn't really are 'factor'

OpenStudy (anonymous):

oh ok thanks